# Epidemiology-informed Graph Neural Network for Heterogeneity-aware Epidemic Forecasting

### v1

@@ IV-A Datasets
  • US-Regions: This dataset includes weekly influenza activity levels (flu patient counts) from 2002 to 2017 for 10 Health and Human Services regions in the United States collected from the Department of Health and Human Services.
  • US-States: This dataset contains weekly influenza disease data (positive cases) from 2010 to 2017 for 49 states in the United States collected from the Center for Disease Control.
The statistics of these three datasets are summarized in Table I. The US-Regions dataset has a higher mean and greater data volatility, the US-States dataset has a lower mean and more concentrated data distribution, and the Japan-Prefectures dataset is between the other two regarding numerical range and data dispersion.
TABLE I: Statistics of the three real-world datasets.
Dataset Size Min Max Mean SD
@@
IV-B Evaluation Metrics
The Root Mean Squared Error (RMSE) measures the difference between predicted and ground truth values after projecting the normalized values into the real range:
$\mathrm{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\hat{y}_{i}-y_{i}\right)^{2}},$ (14)
The PCC is a measure of the linear dependence between two sets of values:
$\mathrm{PCC}=\frac{\sum_{i=1}^{N}\left(\hat{y}_{i}-\overline{\hat{y}}\right)\left(y_{i}-\bar{y}\right)}{\sqrt{\sum_{i=1}^{N}\left(\hat{y}_{i}-\overline{\hat{y}}\right)^{2}}\sqrt{\sum_{i=1}^{N}\left(y_{i}-\bar{y}\right)^{2}}}.$ (15)
where $\hat{y}$ is the predicted values, $y$ is the ground truth values, and $\overline{\hat{y}}$ and $\bar{y}$ are their average values respectively. $N$ represents the number of samples.
IV-C Experimental Settings and Implementation Details
We divide the dataset into training, validation, and test sets in chronological order with a ratio of 60%-20%-20%[29, 20]. We normalized all data based on the training data. The validation set is used for hyperparameter tuning and to prevent overfitting. The test set is used to evaluate the performance of the final model using RMSE and PCC[29]. The smaller the RMSE, the better the performance while the larger the PCC, the better the performance. We used Adam [30] as the optimizer with weight decay 5e-4 during training. The grid search is used to find the best hyperparameter combinations. After the grid search, we set the initial learning rate to 0.001, the early stopping strategy with the patience of 200 epochs, the batch size to 32, and the number of epochs to 1500. All the parameters are initialized using Xavier initialization in HeatGNN. For all models, we set the historical window size $w$ to 20, that is, to use the historical data of the previous 20 weeks as input features. The horizon of advance forecasting $h$ is 2, 3, 4, and 5, that is, to predict the risk of epidemic transmission of the $h$-th week after time step $t$. All experimental results are the average of 5 randomized trials. The experiments are conducted on an Intel Xeon Silver 4214 Processor CPU with 128 GB of 2666 MHz RAM and 4 NVIDIA TITAN RTX 24GB GPUs. Our method implemented in Python 3.8 and PyTorch 1.13.1 trains on GPUs in 30-60 minutes for each forecasting task.
IV-D Comparison Methods
@@
We compared four types of models: traditional statistical models, RNN-based and CNN-based models, STGNN models, and EML models. The traditional statistical models include AutoRegressive (AR), AutoRegressive Moving Average (ARMA), Generalized AutoRegressive (GAR), and Vector AutoRegression (VAR). The RNN-based and CNN-based models include Recurrent Neural Network (RNN)[31], Long Short-Term Memory (LSTM)[32], Gated Recurrent Unit (GRU)[33], RNN-Attn[34], CNNRNN-Res[29], and Long- and Short-term Time-series Network (LSTNet)[35]. The STGNN models include STGCN[27], Multiresolution Graph Neural Networks (MGNN)[15], Temporal Multiresolution Graph Neural Networks (TMGNN)[15], Attention-based Multiresolution Graph Neural Networks (ATMGNN)[16], Cross-location Attention-based Graph Neural Networks (Cola-GNN)[4], and Epi-GNN[12]. The EML model is Epi-Cola-GNN[20].
@@
TABLE II: RMSE and PCC performance of different methods on three datasets with horizon = 2, 3, 4, 5. A boldface indicates the best result for each column, and the second-best is underlined.
Japan-Prefectures US-Regions US-States
RMSE($\downarrow$) 2 3 4 5 2 3 4 5 2 3 4 5
AR 1669.898 2071.535 2312.712 2440.119 625.620 813.453 956.083 1078.889 172.179 210.689 239.205 264.437
ARMA 1661.222 2070.781 2312.378 2438.737 622.393 804.126 939.212 1057.998 172.106 207.789 239.828 262.015
GAR 1484.372 1990.001 2281.875 2427.555 607.232 790.684 934.327 1053.783 161.848 199.566 220.901 247.398
VAR 1577.216 2100.213 2443.453 2566.043 694.516 864.829 1000.032 1112.857 297.669 306.723 318.044 333.312
RNN 1198.314 1551.096 1720.115 1851.500 572.438 753.568 890.801 979.708 158.949 195.903 210.525 224.040
LSTM 1255.470 1555.815 1710.241 1839.544 570.277 743.945 913.847 1048.754 161.029 197.831 216.700 226.667
GRU 1181.230 1498.437 1632.849 1781.807 570.898 741.786 893.440 1024.078 159.298 195.385 217.282 247.858
RNN-Attn 1311.157 1751.819 2056.583 2266.540 566.375 757.980 894.176 999.389 162.821 193.795 219.705 243.932
CNNRNN-Res 1401.394 1889.815 2260.501 2570.203 594.910 782.389 873.559 1006.251 228.275 290.916 308.268 320.370
LSTNet 1441.367 2030.374 2331.407 2385.490 606.540 841.441 1030.450 1134.228 195.949 232.584 255.764 276.306
STGCN 1273.563 1389.538 1498.233 1399.753 726.945 814.875 894.509 965.067 205.060 226.799 246.035 260.379
MGNN 1756.653 1922.766 1822.108 1943.240 1472.420 1415.071 1473.778 1487.338 178.034 189.710 198.840 205.517
TMGNN 1500.051 1471.865 1568.222 1558.965 656.364 755.206 829.295 915.758 178.709 188.653 201.361 190.089
ATMGNN 2436.255 2465.356 2473.451 2464.829 1553.142 1607.391 1691.877 1688.661 307.932 374.719 383.963 355.191
Cola-GNN 1167.715 1416.884 1478.782 1572.566 551.810 681.407 793.208 870.818 147.700 175.602 194.336 223.540
Epi-Cola-GNN 1118.387 2056.956 2282.738 1543.628 546.637 747.352 881.656 871.939 148.455 184.681 204.389 219.132
Epi-GNN 1503.244 1296.059 1284.933 1447.863 563.328 708.990 790.972 881.026 149.515 168.939 184.500 190.005
HeatGNN 1148.850 1260.150 1295.697 1378.207 540.573 693.655 781.513 852.311 142.433 164.788 177.624 186.280
PCC($\uparrow$) 2 3 4 5 2 3 4 5 2 3 4 5
AR 0.743 0.567 0.413 0.303 0.921 0.868 0.829 0.788 0.934 0.903 0.876 0.850
ARMA 0.745 0.567 0.413 0.303 0.920 0.870 0.827 0.786 0.933 0.902 0.873 0.849
GAR 0.803 0.613 0.444 0.318 0.924 0.870 0.823 0.782 0.938 0.906 0.884 0.861
VAR 0.771 0.564 0.369 0.229 0.895 0.839 0.775 0.715 0.779 0.769 0.749 0.730
RNN 0.887 0.810 0.775 0.757 0.934 0.886 0.848 0.819 0.944 0.915 0.902 0.890
LSTM 0.884 0.831 0.818 0.820 0.933 0.890 0.853 0.820 0.943 0.915 0.897 0.892
GRU 0.894 0.840 0.827 0.807 0.933 0.887 0.849 0.817 0.944 0.916 0.897 0.870
RNN-Attn 0.865 0.735 0.589 0.497 0.936 0.889 0.849 0.820 0.943 0.916 0.891 0.869
CNNRNN-Res 0.828 0.653 0.459 0.239 0.925 0.869 0.833 0.777 0.887 0.802 0.769 0.731
LSTNet 0.820 0.586 0.384 0.338 0.934 0.877 0.795 0.733 0.924 0.885 0.848 0.811
STGCN 0.892 0.874 0.852 0.872 0.899 0.866 0.848 0.831 0.901 0.882 0.861 0.838
MGNN 0.688 0.633 0.644 0.507 0.558 0.557 0.557 0.550 0.926 0.914 0.909 0.901
TMGNN 0.827 0.836 0.833 0.822 0.913 0.885 0.870 0.833 0.925 0.916 0.906 0.913
ATMGNN 0.160 0.156 0.162 0.170 0.496 0.481 0.464 0.456 0.737 0.640 0.633 0.698
Cola-GNN 0.906 0.845 0.870 0.834 0.939 0.903 0.872 0.847 0.953 0.936 0.922 0.904
Epi-Cola-GNN 0.912 0.580 0.444 0.859 0.940 0.889 0.851 0.858 0.953 0.922 0.906 0.909
Epi-GNN 0.821 0.911 0.911 0.869 0.937 0.904 0.885 0.858 0.952 0.937 0.925 0.923
HeatGNN 0.917 0.907 0.910 0.884 0.941 0.908 0.885 0.866 0.953 0.939 0.929 0.921
IV-E Overall Comparison
This section addresses RQ1. We evaluate our method in both short-term forecasting (horizon=2, 3) and long-term forecasting (horizon=4, 5) settings. We exclude the case of horizon=1 due to delays in epidemic surveillance data, as extreme short-term forecasting does not provide sufficient time for effective interventions. Table II summarizes the results of all methods under three real-world datasets. While the PCC shows relatively stable ranges and variances between datasets, the RMSE shows significant variability, primarily driven by differences in the variance of the datasets.
@@ IV-E Overall Comparison
Compared with various baseline models, including traditional statistical, RNN-based, CNN-based, STGNN, and EML models, HeatGNN achieves the lowest RMSE and highest PCC across most horizons and datasets. These results demonstrate its capability to capture both spatio-temporal heterogeneity and mechanistic heterogeneity effectively. The high PCC values across datasets and horizons highlight the robustness of HeatGNN in capturing transmission patterns that are closely related to real-world data, making it well-suited for short-term and long-term epidemic forecasting. For instance, in the Japan-Prefectures dataset, HeatGNN achieves a PCC of 0.917 for horizon=2, outperforming the next best model, Epi-Cola-GNN, which scores 0.902. Among the other models, Epi-Cola-GNN and Epi-GNN show the second-best performances. While the forecasting accuracy of Epi-GNN is slightly inferior to HeatGNN in most tasks, it achieves the best performance in a few cases, highlighting its effectiveness and robustness in capturing spatio-temporal dependencies for epidemic forecasting.
@@
As the forecasting horizon increases, HeatGNN maintains competitive performance, while several baseline models show significant error growth. Short-term forecasting performs well due to strong temporal dependencies, as evidenced by good forecasting performance for all models at horizon=2. However, traditional statistical and CNN-based models lose temporal dependency information as the horizon extends, leading to rapid performance declines. In contrast, the RNN-based models maintain relatively robust performance with their memory of temporal dependencies. The improved performance of STGNN models, such as Cola-GNN, STGCN, and TMGNN, highlights the importance of spatio-temporal heterogeneity in epidemic forecasting. Nevertheless, while spatio-temporal dependencies enhance the ability to capture dynamic changes in epidemic transmission, they also introduce added complexity, compromising the robustness of specific models, such as MGNN and ATMGNN. Additionally, incorporating epidemic mechanism models into GNNs improves short-term forecasting accuracy. For instance, Epi-Cola-GNN achieves the second-best performance across all three datasets at horizon=2.
IV-F Ablation Study
This section addresses RQ2. To assess the effect of each component in our framework, we conduct the ablation study by testing the following variants of HeatGNN:
  • HeatGNN w/o PL: Remove the physics loss from HeatGNN removes the constraints on the EIEL module.

### v2

@@ V-A Experimental Settings
  • US-Regions: This dataset includes weekly influenza patient counts from 2002 to 2017 for 10 Health and Human Services regions in the United States collected from the Department of Health and Human Services.
  • US-States: This dataset contains weekly influenza-positive cases from 2010 to 2017 for 49 states in the United States from the Center for Disease Control.
  • Australia-COVID: This dataset obtained from the publicly accessible JHU-CSSE repository comprises daily counts of newly confirmed COVID-19 cases in Australia, encompassing all six states and two territories, spanning the period from January 27, 2020, to August 4, 2021.
The statistics of these four datasets are shown in Table I. We use Root Mean Squared Error (RMSE) and PCC to measure the difference between predicted and ground truth values:
$\begin{gathered}\mathrm{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\hat{y}_{i}-y_{i}\right)^{2}},\\ \mathrm{PCC}=\frac{\sum_{i=1}^{N}\left(\hat{y}_{i}-\overline{\hat{y}}\right)\left(y_{i}-\bar{y}\right)}{\sqrt{\sum_{i=1}^{N}\left(\hat{y}_{i}-\overline{\hat{y}}\right)^{2}}\sqrt{\sum_{i=1}^{N}\left(y_{i}-\bar{y}\right)^{2}}},\end{gathered}$ (14)
where $\hat{y}$ is the predicted values, $y$ is the ground truth values, and $\overline{\hat{y}}$ and $\bar{y}$ are their average in $N$ samples respectively.
TABLE I: Statistics of the four real-world datasets.
Dataset Size Min Max Mean Standard Deviation
@@
Australia-COVID $8\times 556$ 0 9987 539 1532
@@
Comparison Methods. We compare four types of models: The traditional models include SIR, AutoRegressive (AR), AutoRegressive Moving Average (ARMA), Generalized AutoRegressive (GAR), and Vector AutoRegression (VAR). The Recurrent Neural Network (RNN)-based and Convolutional Neural Network (CNN)-based models include RNN [werbos1990backpropagation], Long Short-Term Memory (LSTM) [graves2012long], Gated Recurrent Unit (GRU) [dey2017gate], RNN-Attn [cheng2016long], CNNRNN-Res [wu2018deep], and Long- and Short-term Time-series Network (LSTNet) [lai2018modeling]. The STGNN models include STGCN [yu2017spatio], Multiresolution Graph Neural Networks (MGNN) [hy2022temporal], Temporal Multiresolution Graph Neural Networks (TMGNN) [hy2022temporal], Cross-location Attention-based Graph Neural Networks (Cola-GNN) [deng2020cola], and Epi-GNN [xie2022epignn]. The epidemiological machine learning model is Epi-Cola-GNN [liu2023epidemiology].
Implementation Details. We divide the dataset into training, validation, and test sets in chronological order with a ratio of 60%-20%-20% [wu2018deep, liu2023epidemiology]. We normalize all data based on the training data. The validation set is used for hyperparameter tuning and to prevent overfitting. The test set is used to evaluate the performance of the final model using RMSE and PCC [wu2018deep]. The smaller the RMSE, the better the performance while the larger the PCC, the better the performance. We used Adam [kingma2014adam] as the optimizer with weight decay 5e-4 during training. The grid search is used to find the best hyperparameter combinations. After the grid search, we set the initial learning rate to 0.001, the early stopping strategy with the patience of 200 epochs, the batch size to 32, and the number of epochs to 1500. All the parameters are initialized using Xavier initialization in HeatGNN. For all models, we set the historical window size $w$ to 20, that is, to use the historical data of the previous 20 weeks as input features. The horizon of advance forecasting $h$ is 2, 5, 7, and 12, that is, to predict the infection number of the $h$-th week after time step $t$. All experimental results are the average of 5 randomized trials. The experiments are conducted on an Intel Xeon Silver 4214 Processor CPU with 128 GB of 2666 MHz RAM and 4 NVIDIA TITAN RTX 24GB GPUs. Our method implemented in Python 3.8 and PyTorch 1.13.1 trains on GPUs in 30-60 minutes for each forecasting task.
V-B Prior Experiment
To explore the heterogeneity of the spread of the epidemic in different regions, we design a prior experiment. We use the time-varying SIR model (Equation 3) to estimate the changes in the number of infections and the changes in its time-varying parameters in each location in Japan-Prefectures in the first 54 weeks (1 year). Based on previous studies, we use the simulated annealing algorithm to estimate the time-varying parameters (infection rate $\beta$ and recovery rate $\gamma$) in the time-varying SIR model. Note that each location is modeled and estimated separately. The experimental results show that the PCC between the true value and the estimated value of each location is basically above 0.9 (Table II). These results show that the fitting results are accurate and effective (Figure 3(a)). After ensuring that the estimation results are reliable, we observe the heterogeneity of the time-varying parameters between different locations. We found that even if two locations have similar spatio-temporal transmission trends (e.g., infection number), there are differences in the trends of their internal mechanism transmission (e.g., infection rate and recovery rate). Some changes in the internal mechanism transmission are earlier than the easily observable change in spatio-temporal transmission (Figure 1). For example, Figure 3 shows the change in the infection number at one location and the change in the time-varying parameters (infection rate $\beta$ and recovery rate $\gamma$) estimated based on the time-varying SIR model. From week 10 to week 15, the estimated infection rate increased before the infection number increased, accompanied by a decrease in the recovery rate.
Moreover, estimating time-varying parameters for all locations is computationally intensive. At each location, the two parameters to be estimated for the time-varying SIR model at a given moment are the infection rate and the recovery rate. If the parameters are estimated for all locations, with a time period $T$ and $N$ locations, the total number of estimated parameters is $2\cdot T\cdot N$. Typically, the estimation of epidemic model parameters relies on metaheuristic algorithms, which present a significant challenge due to their computational complexity [zhan2021identifying, zelenkov2023analysis]. This shows that it is unrealistic to use metaheuristic algorithms in deep learning models to estimate parameters for epidemic forecasting.
TABLE II: Statistics of the fitting of the prior experiment.
Mean Max Min Standard Deviation
PCC 0.9267 0.9983 0.8145 0.0462
RMSE 704.33 3400.30 18.039 825.43
Figure 3: Example of prior experiment. (a) is the estimation curve of infection number. (b) is the change of estimation parameter.
V-C Overall Comparison (RQ1)
@@
TABLE III: Performance results on four datasets with horizon = 2, 5, 7, 12. A boldface indicates the best result, and the second-best is underlined. RMSE values are scaled by $\times 10^{3}$ to save space.
Method Japan-Prefectures US-Region
RMSE ($\downarrow$) PCC ($\uparrow$) RMSE ($\downarrow$) PCC ($\uparrow$)
2 5 7 12 2 5 7 12 2 5 7 12 2 5 7 12
SIR 78.173 217.438 432.737 1991.195 0.071 -0.077 -0.124 -0.150 71.304 125.453 182.120 446.533 0.533 0.422 0.348 0.183
AR 1.670 2.440 2.561 2.485 0.743 0.303 0.154 0.491 0.626 1.079 1.280 1.496 0.921 0.788 0.705 0.583
ARMA 1.661 2.439 2.559 2.482 0.745 0.303 0.154 0.495 0.622 1.058 1.263 1.484 0.920 0.786 0.699 0.576
GAR 1.484 2.428 2.561 2.530 0.803 0.318 0.152 0.413 0.607 1.054 1.288 1.568 0.924 0.782 0.690 0.520
VAR 1.577 2.566 2.530 2.504 0.771 0.229 0.170 0.359 0.695 1.113 1.234 1.380 0.895 0.715 0.636 0.529
RNN 1.198 1.852 2.193 1.906 0.887 0.757 0.593 0.726 0.572 0.980 1.154 1.513 0.934 0.819 0.756 0.537
LSTM 1.255 1.840 2.235 1.952 0.884 0.820 0.606 0.723 0.570 1.049 1.230 1.531 0.933 0.820 0.754 0.580
GRU 1.181 1.782 2.226 1.922 0.894 0.807 0.599 0.725 0.571 1.024 1.205 1.521 0.933 0.817 0.753 0.527
RNN-Attn 1.311 2.267 2.133 1.933 0.865 0.497 0.662 0.715 0.566 0.999 1.211 1.521 0.936 0.820 0.734 0.567
CNNRNN-Res 1.401 2.570 2.603 2.067 0.828 0.239 0.076 0.654 0.595 1.006 1.153 1.316 0.925 0.777 0.687 0.600
LSTNet 1.441 2.385 2.477 2.391 0.820 0.338 0.267 0.438 0.607 1.134 1.121 1.149 0.934 0.733 0.748 0.682
STGCN 1.274 1.400 1.478 1.812 0.892 0.872 0.853 0.761 0.727 0.965 1.017 1.116 0.899 0.831 0.817 0.733
MGNN 1.757 1.943 2.048 2.655 0.688 0.507 0.400 0.133 1.472 1.487 1.305 1.538 0.558 0.550 0.549 0.536
TMGNN 1.500 1.559 1.977 1.839 0.827 0.822 0.721 0.777 0.656 0.916 1.043 1.228 0.913 0.833 0.808 0.796
Cola-GNN 1.168 1.573 1.758 1.690 0.906 0.834 0.787 0.804 0.552 0.871 1.054 1.011 0.939 0.847 0.805 0.780
Epi-Cola-GNN 1.118 1.544 2.549 2.164 0.912 0.859 0.179 0.679 0.547 0.872 1.160 1.365 0.940 0.858 0.751 0.648
Epi-GNN 1.503 1.448 1.738 1.695 0.821 0.869 0.768 0.786 0.563 0.881 0.938 1.025 0.937 0.858 0.849 0.821
HeatGNN 1.149 1.378 1.735 1.685 0.917 0.884 0.780 0.773 0.541 0.852 0.922 1.024 0.941 0.866 0.856 0.824
Method US-State Australia-COVID
RMSE ($\downarrow$) PCC ($\uparrow$) RMSE ($\downarrow$) PCC ($\uparrow$)
2 5 7 12 2 5 7 12 2 5 7 12 2 5 7 12
SIR 17.872 31.446 45.436 110.536 0.656 0.584 0.537 0.431 74.009 112.709 144.578 231.556 0.996 0.995 0.993 0.989
AR 0.172 0.264 0.300 0.341 0.934 0.850 0.795 0.717 0.509 0.540 0.556 0.587 0.990 0.989 0.988 0.986
ARMA 0.172 0.262 0.298 0.340 0.933 0.849 0.796 0.718 0.505 0.527 0.541 0.564 0.989 0.988 0.987 0.986
GAR 0.162 0.247 0.292 0.355 0.938 0.861 0.805 0.708 1.343 1.394 1.430 1.519 0.733 0.727 0.723 0.700
VAR 0.298 0.333 0.397 0.422 0.779 0.730 0.685 0.581 0.581 0.592 0.607 0.619 0.984 0.982 0.981 0.978
RNN 0.159 0.224 0.255 0.329 0.944 0.890 0.861 0.747 0.581 0.648 0.708 0.898 0.992 0.991 0.991 0.990
LSTM 0.161 0.227 0.263 0.330 0.943 0.892 0.853 0.749 1.122 1.178 1.230 1.393 0.985 0.984 0.984 0.982
GRU 0.159 0.248 0.285 0.335 0.944 0.870 0.822 0.742 1.065 1.127 1.177 1.322 0.995 0.994 0.993 0.991
RNN-Attn 0.163 0.244 0.274 0.342 0.943 0.869 0.823 0.711 1.054 1.039 1.026 1.166 0.862 0.867 0.868 0.717
CNNRNN-Res 0.228 0.320 0.360 0.280 0.887 0.731 0.649 0.808 1.484 1.488 1.491 1.476 0.959 0.962 0.961 0.964
LSTNet 0.196 0.276 0.288 0.312 0.924 0.811 0.785 0.794 0.595 0.614 0.645 0.712 0.990 0.989 0.988 0.984
STGCN 0.205 0.260 0.280 0.296 0.901 0.838 0.808 0.804 0.696 0.649 0.706 0.721 0.960 0.961 0.964 0.971
MGNN 0.178 0.206 0.259 0.288 0.926 0.901 0.870 0.851 0.664 0.769 0.848 0.857 0.826 0.823 0.820 0.818
TMGNN 0.179 0.190 0.438 0.447 0.925 0.913 0.549 0.545 0.803 0.846 0.870 0.916 0.829 0.828 0.827 0.824
Cola-GNN 0.148 0.224 0.228 0.241 0.953 0.904 0.898 0.893 0.368 0.406 0.463 0.533 0.993 0.994 0.990 0.988
Epi-Cola-GNN 0.148 0.219 0.268 0.303 0.953 0.909 0.850 0.819 0.803 0.854 0.912 1.124 0.974 0.975 0.957 0.970
Epi-GNN 0.150 0.190 0.205 0.251 0.952 0.923 0.909 0.881 0.325 0.381 0.416 0.491 0.995 0.993 0.992 0.989
HeatGNN 0.142 0.186 0.200 0.240 0.953 0.921 0.911 0.870 0.315 0.334 0.411 0.466 0.993 0.993 0.991 0.986
@@
We evaluate our method in both short-term forecasting (horizon=2, 5) and long-term forecasting (horizon=7, 12) settings. Table III summarizes the results of all methods under four real-world datasets, where we scale the RMSE to save space. Compared with various baseline models, including traditional, RNN-based, CNN-based, STGNN, and EML models, HeatGNN achieves the lowest RMSE and highest PCC across most horizons and datasets. These results demonstrate its capability to capture both spatio-temporal heterogeneity and mechanistic heterogeneity effectively. The high PCC values across datasets and horizons highlight the robustness of HeatGNN in capturing transmission patterns that are closely related to real-world data, making it well-suited for short-term and long-term epidemic forecasting. Among the other models, Epi-Cola-GNN and Epi-GNN show the second-best performances. Figure 4 shows that HeatGNN more accurately fits the ground truth and follows the epidemic transmission trend.
Figure 4: Visualization of prediction and ground truth values under horizon=5 in US-Regions.
@@
As the horizon increases, HeatGNN maintains competitive performance, while several models show significant error growth. Short-term forecasting performs well due to strong temporal dependencies, as evidenced by good forecasting performance for all models at horizon=2. However, traditional and CNN-based models lose temporal dependency information as the horizon extends, leading to performance declines. In contrast, the RNN-based models maintain relatively robust performance with their memory of temporal dependencies. The improved performance of STGNN models, such as Cola-GNN, STGCN, and TMGNN, highlights the importance of spatio-temporal heterogeneity in forecasting. However, while spatio-temporal dependencies enhance the ability to capture dynamic changes in epidemic transmission, they also introduce added complexity, compromising the robustness of specific models, such as MGNN and TMGNN. Additionally, incorporating epidemic mechanism models into GNNs improves short-term forecasting accuracy. For instance, Epi-Cola-GNN achieves the second-best performance across three datasets at horizon=2.
V-D Ablation Study (RQ2)